# Ring Deconvolution Microscopy: An Exact Solution for Spatially-Varying Aberration Correction

### v3

@@ 6 Data Availability
The data used in all of the imaging experiments (Miniscope, Multicolor fluorescence, and multimode fiber) is publicly available on Box (https://berkeley.box.com/s/zmsjjgmquwq2roh4d9qthcnv3rhwuidn). Additional experimental data from the multimode fiber system can be requested from [15](https://opg.optica.org/boe/fulltext.cfm?uri=boe-11-8-4759&id=433935). The datasets used to train and fine-tune DeepRD, and to evaluate the quantitative performance of the methods are also hosted on Box (https://berkeley.box.com/s/vv3g6avhrr9agijmlj3b1153oo7x9gao). These datasets were sourced from the CARE dataset [21] (https://publications.mpi-cbg.de/publications-sites/7207/) and the Div2k dataset [37] (https://data.vision.ee.ethz.ch/cvl/DIV2K/). The high resolution pretraining dataset, due to its large memory useage, will be made available upon request.
@@ 7 Inclusion and Ethics
This study included contributions from local researchers at UC Berkeley, UC San Francisco, and Nikon Imaging Center at Harvard Medical School. Data was acquired by local researchers at each location. These researchers are either included as authors or acknowledged below.
@@ Acknowledgements
We would like to thank Ana Lyons and Saul Kato’s lab for providing the tardigrades for imaging. We would like to thank the authors of [15] for giving us access to the micro-endoscopy data. We would also like to thank the Nikon Imaging Center at Harvard Medical School for the imaging support. We would like to acknowledge Neerja Aggarwal for help and access to 3D printing. A.K. was funded by the Berkeley Fellowship for Graduate Study. A.N.A. was supported by the Berkeley Fellowship for Graduate Study and the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 1752814. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. S.Y. is supported by Moore Foundation data-driven discovery. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0521. This publication has been made possible in part by CZI grant DAF2021-225666 and grant DOI https://doi.org/10.37921/192752jrgbn from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation (funder DOI 10.13039/100014989).
Appendix A Interpretability of DeepRD
One key property of any deep learning model is its interpretability. How well we understand what the network is doing to produce its predictions corresponds to how confident we are in its performance. One such mark of interpretability when working with encoder-decoder frameworks, is whether the latent space, i.e., the space of encodings, is disentagled, meaning each coordinate of the encoding has a distinctive meaning. In the case of DeepRD, the latent space is the space of Seidel coefficients. In the ideal case we expect the network to learn about the individual effects of each coordinate (each aberration coefficient) it takes in as input; say we only increase the coefficient value for radial distortion, then we expect the network to further unwarp the predicted image correspondingly.
This property indeed emerges in practice. We observe this by deblurring images of insect cornea taken by the Miniscope using a sweep over each aberration coefficient independently. For each coefficient, we set all other coefficients to 0 and sweep the chosen coefficient from 0 to 3 waves. What we observe is that the network only corrects for the chosen aberration in increasing amounts while leaving the other characteristics of the image alone. For example, when we sweep distortion, the image remains blurry but radially warps in the opposite direction compared to if it were distorted. Results for each aberration coefficient are shown in video form here. Such results confirms our hypothesis that DeepRD is indeed learning to perform ring deconvolution given a systems Seidel coefficients, and does not rely on knowledge of the specific image distribution. This further aids our confidence in its ability to generalize.
Appendix B Seidel Coefficients
@@
Here, we give a brief background on Seidel coefficients. This is not a complete treatment and we encourage an interested reader to refer to Voelz et al.[50] for an overview, and the standard Born et al.[34] for a fully rigorous derivation.
@@
Consider an rotationally symmetric—and consequently LRI—imaging system. It is common to consolidate the system aberrations into a single function and apply it at the exit pupil plane of the system. This complex-valued function, known as the generalized pupil function $p$ is composed of a binary-valued amplitude distribution (set by the shape of the pupil) and a phase distribution $w$, which quantifies the deviation of the pupil wavefront from the ideal spherical shape necessary for diffraction-limited imaging. Since LRI systems are generally spatially-varying (albeit only radially), $w$ becomes a function of the distance of the source from the optical axis $r$. Letting $(s,t)$ be the pupil plane coordinates, we can write $p$ as
@@
For rotationally symmetric systems, as in our case, it is possible to expand $w$ as an infinite power series (see the aberration function in [34]). Usually only the fourth-order terms of this series are used, which yields the following 2D polynomial of the pupil plane coordinates,
@@
Appendix C A Fourier analysis of LRI systems
@@
Now our notion of resolution, assuming the ability to perfectly de-mix, is radially dependent and, at radius $r$, is proportional to the angular arc length at $\rho=r$ of the $r^{th}$ PSF. The LRI filtering interpretation offers a more realistic understanding of how an imaging system filters an object and opens the door for a host of new imaging techniques which optimize for key features, such as resolution, under the LRI model.
Appendix D Additional Related Work
The idea of linear space-invariance (LSI)—or rather its 1D predecessor time invariance—is difficult to trace historically, but the first mention of it that we were able to find was by Richard Hamming in 1934 [56], followed by the far more popular definition provided in [57]. The concept also has a close connection to the convolution integral—which had existed separately as far back as the 1800s [58]—through the famous Convolution Theorem. Much of the foundational mathematics of convolution and its frequency domain interpretation comes from the work of Laplace and Fourier; these techniques enabled the application of linear time (space)-invariant systems to communications and, of course, imaging. The practical use of LSI systems for imaging analysis arose from Bell labs during the emergence of digital signal processing for communication in the 1960s and 1970s.
Image deblurring has its roots much later; it was only in 1949 when Norbert Wiener published the Wiener Filter for linear time-invariant deconvolution [59]. While the original technique was developed in 1940 for time series data in order to track enemy planes in World War II, it was soon adapted for image deblurring. With the rise of digital computing in the following decades came an explosion of image deconvolution techniques which are commonly employed today such as Richardson-Lucy deconvolution [60, 61]. It is interesting to note that some of these algorithms have existed for nearly a century and yet there is no clear “best” method.

### v5

@@ 6 Data Availability
The data used in all of the imaging experiments (Miniscope, multicolor fluorescence, multimode fiber, and light-sheet) is publicly available on Box (https://berkeley.box.com/s/zmsjjgmquwq2roh4d9qthcnv3rhwuidn). Additional experimental data from the multimode fiber system can be requested from [Turcotte:20](https://opg.optica.org/boe/fulltext.cfm?uri=boe-11-8-4759&id=433935). The datasets used to train and fine-tune DeepRD, and to evaluate the quantitative performance of the methods are also hosted on Box (https://berkeley.box.com/s/vv3g6avhrr9agijmlj3b1153oo7x9gao). These datasets were sourced from the CARE dataset [weigert2018content] (https://publications.mpi-cbg.de/publications-sites/7207/) and the Div2k dataset [agustsson2017ntire] (https://data.vision.ee.ethz.ch/cvl/DIV2K/). The high resolution pretraining dataset, due to its large memory useage, will be made available upon request.
@@ 7 Inclusion and Ethics
This study included contributions from local researchers at UC Berkeley, UC San Francisco, UT Southwestern, and Nikon Imaging Center at Harvard Medical School. Data was acquired by local researchers at each location. These researchers are either included as authors or acknowledged below.
@@ Acknowledgements
We would like to thank the following organizations and individuals: Ana Lyons and Saul Kato’s lab for providing the tardigrades for imaging, the authors of [Turcotte:20] for giving us access to the micro-endoscopy data, the Nikon Imaging Center at Harvard Medical School for their imaging support, Neerja Aggarwal at UC Berkeley for help and access to 3D printing, Gabriel Muhire Gihana from Danuser lab at UTSW for providing SU8686 cells, and Diptodip Deb from Janelia Research Campus in Virginia for running preliminary multi-GPU experiments.
@@
A.K. was funded by the Berkeley Fellowship for Graduate Study. A.N.A. was supported by the Berkeley Fellowship for Graduate Study and the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 1752814. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. S.Y. is supported by Moore Foundation data-driven discovery. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-22-1-0521. This publication has been made possible in part by CZI grant DAF2021-225666 and grant DOI https://doi.org/10.37921/192752jrgbn from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation (funder DOI 10.13039/100014989).
[…]
Appendix A Additional comparisons
Two other commonly used deconvolution methods are Gaussian kernel deconvolution and modal decomposition. Gaussian kernel deconvolution involves fitting a two-dimensional Gaussian distribution to an experimental PSF and using the fitted kernel for deconvolution. While common, this method assumes spatial-invariance and cannot correct spatially-varying aberrations. On the other hand, modal decomposition is a spatially-varying deconvolution technique which models the spatially-varying image formation as a weighted sum of convolutions with space-invariant kernels. These kernels and their corresponding weights can be estimated from a uniformly spaced set of PSFs via singular value decomposition. We note that this type of calibration generally cannot be done single-shot and requires multiple acquisitions with a high precision motion stage. Since the multimode fiber experiment has such calibration information, we use it as a test bed for these additional techniques. Figure 8 shows these methods on a grid of beads taken with the same multimode fiber system used in [Turcotte:20].
Ring deconvolution exhibits the best performance among these methods. Gaussian kernel deconvolution suffers from a lack of detail in its PSF compared to the experimental one. Meanwhile modal decomposition, despite using a grid of 21 by 21 PSFs, is unable to properly interpolate the PSFs in areas between the calibration points and provides erratic performance throughout the FoV. It is also worth noting that that modal decomposition could be combined with our Seidel fitting procedure to further improve its accessibility; this exploration however is beyond the scope of this work and is left for future exploration.
Figure 8: Comparisons of various deblurring methods. An evenly spaced grid of 13 x 13 point sources are imaged with the multimode fiber micro-endoscope from [Turcotte:20] and deblurred using a variety of methods. a) Results from each method. Top row (left to right) is the raw measurement, deconvolution with an experimental PSF and deconvolution with a Gaussian kernel fitted to the PSF. Bottom row (left to right) is deblurring via modal decomposition [yanny2020miniscope3d, flicker2005anisoplanatic], SVRL [Turcotte:20], and ring deconvolution. As seen in b), the spatially-varying methods (bottom row) are superior to the deconvolution methods (top row) due to the substantial spatial variation in the PSF. Among the spatially-varying methods, ring deconvolution produces the smallest, most consistent beads. The quantitative results in c) also show that ring deconvolution has the smallest average bead radius with the least variation. Moreover, unlike the other methods, its performance does not degrade on beads far from the center of the FoV.
Appendix B Quantifying the performance of SDM
Figure 9: Quantifying Sheet deconvolution microscopy. a) Error evaluation of PSF fitting models. N=6 image stacks of randomly-scattered point sources are acquired with a light-sheet microscope and fit using either with the modified Gibson-Lanni model, by interpolating the measured PSFs, or by using the (denoised) center PSF and assuming shift-invariance. The PSNRs between the generated PSFs and the measured PSFs from the remaining calibration image stacks is computed for all three PSF models. These PSNRs are stratified by regions split along the light-sheet spread direction; center is where the light-sheet is the thinnest, while edge is where the light-sheet is the most spread. The resulting average PSNRs and their standard deviations in each region are plotted as a function of region. b) Forward model computation time for different data sizes. The runtimes of three different light-sheet forward models are plotted as a function of data size; true blur (manually superimposing the shift-varying PSFs at every point in the 3D volumes), sheet convolution, and standard 3D convolution. Times are acquired by processing randomly-generated 3D volumes and are averaged over 10 trials. c) Quantitative evaluation of sheet deconvolution. N=20 3D volumes of randomly sized and oriented ellipsoids are generated. These objects are blurred via true blur using PSFs generated from our modified Gibson-Lanni model, and are corrupted with Poisson noise (SNR 15) to simulate light-sheet measurements. The simulated measurements are deblurred with sheet deconvolution, iterative gradient-based deconvolution with total variation regularization, Richardson-Lucy deconvolution, and Wiener deconvolution. The process is repeated with increasing values of the $\beta$ spread parameter (see 15), beginning with a system that is shift-invariant ranging to a system which is highly shift-variant due to rapid light-sheet spread. The average PSNR over all 20 objects for each method is plotted as a function of the spread parameter.
The experimental evaluation of sheet deconvolution microscopy (SDM) (Sec. 2.4) highlights its resolution improvement and computational efficiency advantages. We further support these results by quantifying each part of the SDM pipeline. Figure 9 details these quantitative experiments. First, quantify how well our light-sheet PSF fitting models, modified Gibson-Lanni and interpolation, fit experimental PSFs and compare them to the standard shift-invariant model. To that end, we fit these models to experimental PSFs acquired from a light-sheet microscope and compute the PSNR between generated PSFs and measured PSFs. We plot these PSNRs as a function of the PSF’s location along the light-sheet spread, from the center, where the light-sheet is the thinnest, to the edges, where the light sheet is the most spread. Both of our models account for the shift-variance due to the light-sheet spread and consequently maintain accurate fitting across the entire FoV. In contrast, the standard shift-invariant model degrades away from the center, as expected. The interpolation model is higher variance than the modified Gibson-Lanni model; this is because the interpolation model is exact at the locations of the PSFs provided in the calibration image and worse in-between those PSFs, whereas the Gibson-Lanni fitting finds parameters that produce PSFs that fit well everywhere in the FoV, on average.
Next, we measure the runtime of the corresponding forward models, sheet convolution and 3D convolution, on different data sizes and find that, while sheet convolution is slower than 3D convolution, it is much faster than computing the true blur, i.e., manually superimposing the PSF at each point in the 3D volume. The scaling of runtime with data size of sheet deconvolution and 3D deconvolution can be inferred from these results but ultimately also depends on the sample and noise properties.
Finally, we quantify the accuracy of sheet deconvolution against standard deconvolution techniques, total variation (TV) iterative deconvolution, Richardson-Lucy deconvolution [Richardson, Lucy], and Wiener deconvolution [Orieux:10]. To do this we simulate phantom 3D samples (randomly sized and oriented ellipsoids) and simulate a light-sheet measurement using the true blur model with Poisson noise. We then deblur the simulated measurement with each method and compute the PSNR of the result with the ground truth sample. We repeat this process for systems with increasing light-sheet spread (increasing spatial variance). When the system is shift-invariant, sheet deconvolution performs as well as the best deconvolution technique, TV regularized iterative deconvolution. As the system becomes increasingly spatially-variant, the performance of sheet deconvolution remains relatively constant, while the performance of all the deconvolution methods degrades severely.
Appendix C Simulation with Poisson noise
The quantitative experiments from Fig. 5 are rerun with Poisson noise instead of Gaussian noise. There are no notable differences in performances, with the same trends occurring as before: ring deconvolution is the best performing method, followed by DeepRD, the baseline U-Net, and finally standard deconvolution. Both deep learning models were updated with a small amount of training (100100~{}100100 iterations) on Poisson-noised images from the training set. Ring deconvolution and standard deconvolution were run with additional total variation regularization to account for the Poisson noise. We found that an SNR range of 15-30 most accurately reflected experimental data and thus ran all simulations with the worst-case SNR of 15.
Figure 10: Quantitative performance under Poisson noise. a) A Poisson-noised image (SNR 15) of randomly placed PSFs is used to estimate the underlying Seidel coefficients, which can then be used to generate PSFs at any location. e) Seidel fitting error is plotted as a function of iteration in the optimization algorithm to demonstrate convergence. Each purple line is a different trial of n=50 trials with a randomly sampled set of underlying Seidel coefficients. The red dashed line is the per-iteration median. b) Mean squared error of the fitted Seidel coefficients plotted as a function of SNR of the calibration image. The average coefficient error is plotted along with the variance (error bar) over n=50 random trials. Some example calibration PSFs are shown. Even under severe noise, the Seidel fit is still accurate. c) Accuracy (PSNR) vs runtime of each method (averaged over n=28 true blurred images using unseen coefficients), with the number of model parameters (written below each circle) determining the size of the circle. Noise is Poisson distributed with SNR 15.
Appendix D Seidel Coefficients
Figure 11: Approximation of higher-order Seidel coefficients with primary Seidel coefficients. a) An example trial: a random set of 13 Seidel coefficients are selected and used to generate a noisy calibration image. Next, only 5 primary Seidel coefficients are fit to the calibration image and are used to generate estimated PSFs. An error map between the high-order and fitted PSFs is shown at the bottom; some of the high-frequency features are missed by the lower-order approximation. b) Error in the fit between the estimated primary coefficients and true primary coefficients in the presence of higher-order coefficients. 50 trials are plotted with the red curve indicating the median convergence. The correct primary coefficients are estimated a majority of the time. We hypothesize that the small bump in the plot where the error slightly increases before convergence is an inflation of the primary coefficients to better explain the higher-order PSFs. c) Images blurred with higher-order PSFs are deblurred with RDM using lower order approximations of the PSFs and the PSNRs are recorded. The experiment is repeated with increasing degrees of higher-order blur, and thus increasing discrepancy from the lower-order approximations. The average PSNR and its standard deviation over 50 trials for ring deconvolution and DeepRD are plotted as a function of the number of higher-order coefficients used in the blur, from 0 (baseline case) to 8 (all 13 coefficients). While the average PSNRs for ring deconvolution seem to drop slightly from the baseline, all orders for both methods have PSNRs within a standard deviation of the baseline error, suggesting the presence of higher-order coefficients are not enough to drastically affect RDM. For all experiments, noise was Poisson with SNR 30 and the coefficients were always normalized to unit norm.
@@ Appendix D Seidel Coefficients
Here, we give a brief background on Seidel coefficients. This is not a complete treatment and we encourage an interested reader to refer to Voelz et al.[voelz2011computational] for an overview, and the standard Born et al.[born2013principles] for a fully rigorous derivation.
@@
Consider a rotationally-symmetric—and consequently LRI—imaging system. It is common to consolidate the system aberrations into a single function and apply it at the exit pupil plane of the system. This complex-valued function, known as the generalized pupil function $p$ is composed of a binary-valued amplitude distribution (set by the shape of the pupil) and a phase distribution $w$, which quantifies the deviation of the pupil wavefront from the ideal spherical shape necessary for diffraction-limited imaging. Since LRI systems are generally spatially-varying (albeit only radially), $w$ becomes a function of the distance of the source from the optical axis $r$. Letting $(s,t)$ be the pupil plane coordinates, we can write $p$ as
@@
For rotationally-symmetric systems, as in our case, it is possible to expand $w$ as an infinite power series (see the aberration function in [born2013principles]). Usually only the 4th order terms of this series are used, which yields the following 2D polynomial of the pupil plane coordinates,
@@
D.1 Higher-order coefficients
The Seidel series is an infinite power series expansion and thus has an infinite number of coefficients. While the aforementioned primary coefficients are dominant in practical optics—particularly for spherical lenses—it is worth considering the impact of higher-order coefficients, especially in the context of our Seidel fitting and ring convolution/deconvolution algorithms. To that end, we include a simulated experiment that tests our methods in the additional presence of 6th order Seidel aberration coefficients. In the experiment (shown in Fig. 11), a randomly-generated set of 13 (5 primary and 8 6th order) coefficients with unit norm are used to generate PSFs, which are then used to create a calibration image and blur a test image. The calibration image is noised and used to fit only the 5 primary coefficients. Finally, the fitted primary coefficients are used to deblur the blurry test image via ring deconvolution. Despite a noticeable difference in the higher-order PSFs, the fitting procedure almost always was able to converge to the correct primary coefficients, albeit with a slightly inflated values—likely accounting for the effects of the higher-order terms. The average deblurring error for ring deconvolution drops slightly compared to the baseline case (0 high-order coefficients), but is within a standard deviation of the baseline error. DeepRD’s error also remains virtually unchanged and appears even more robust to the addition of high-order coefficients. Thus even in the presence of higher-order aberrations, RDM calibrated with only the primary Seidel coefficients produces accurate deblurring. As a final note, though the primary coefficients are largely sufficient, we do additionally include the option to fit the higher-order terms in our codebase.
Appendix E Interpretability of DeepRD
One desirable property of DeepRD is its interpretability. Because the DeepRD latent space lives in the space of Seidel coefficients, we can easily see why our image reconstruction looks the way it does. Moreover, if our reconstruction is deficient, we can manually adjust the latent space, thereby changing the Seidel coefficients and improving the reconstruction. In the literature on encoder-decoder frameworks, this property is called disentanglement, that is, each element in the latent space has its own distinct and interpretable meaning. For example, if we only increase the coefficient value for radial distortion, then we expect the network to further unwarp the predicted image correspondingly.
This disentanglement property emerges in practice. We observe this by deblurring images of insect cornea taken by the Miniscope using a sweep over each aberration coefficient independently. For each coefficient, we set all other coefficients to 0 and sweep the chosen coefficient from 0 to 3 waves. What we observe is that the network only corrects for the chosen aberration in increasing amounts, while leaving the other characteristics of the image alone. For example, when we sweep distortion, the image remains blurry, but radially warps in the opposite direction compared to if it were distorted. Results for each aberration coefficient are shown in video form here. Such results confirm our hypothesis that DeepRD is indeed learning to perform ring deconvolution given a system’s Seidel coefficients, and does not rely on knowledge of the specific image distribution. This further aids our confidence in its ability to generalize.
Appendix F A Fourier analysis of LRI systems
@@
Now our notion of resolution, assuming the ability to perfectly de-mix, is radially dependent and, at radius $r$, is proportional to the angular arc length at $\rho=r$ of the $r^{th}$ PSF. The LRI filtering interpretation offers a more realistic understanding of how an imaging system filters an object and opens the door for a host of new imaging techniques which optimize for key features, such as resolution, under the LRI model. An analogous analysis for systems with other forms of symmetry— like the linear symmetry present in light-sheet microscopy—will lead to similar conclusions.
Appendix G Additional Related Work
The idea of linear space-invariance (LSI)—or rather its 1D predecessor time invariance—is difficult to trace historically, but the first mention of it that we were able to find was by Richard Hamming in 1934 [hamming], followed by the far more popular definition provided in [oppenheim1975digital]. The concept also has a close connection to the convolution integral—which had existed separately as far back as the 1800s [dominguez_2015]—through the famous Convolution Theorem. Much of the foundational mathematics of convolution and its frequency domain interpretation comes from the work of Laplace and Fourier; these techniques enabled the application of linear time (space)-invariant systems to communications and, of course, imaging. The practical use of LSI systems for imaging analysis arose from Bell labs during the emergence of digital signal processing for communication in the 1960s and 1970s.
Image deblurring has its roots much later; it was only in 1949 when Norbert Wiener published the Wiener Filter for linear time-invariant deconvolution [wiener1964extrapolation]. While the original technique was developed in 1940 for time series data in order to track enemy planes in World War II, it was soon adapted for image deblurring. With the rise of digital computing in the following decades came an explosion of image deconvolution techniques which are commonly employed today such as Richardson-Lucy deconvolution [Richardson, Lucy]. It is interesting to note that some of these algorithms have existed for nearly a century and yet there is no clear “best” method.